OpenStudy (anonymous):

Please help check my answers! Let E and F be events such that P(E)=0.45 and P(F)=0.17. Find: a. P(E^c) -> I got 0.55 b. P(not F) -> I got 0.38 c. P(E or F) assuming E and F are mutually exclusive -> I got 0.62 d. P(E) or (F) assuming independent -> I got 61.9235

jimthompson5910 (jim_thompson5910):

P(not F) should be 0.83 since P(not F) + P(F) = 1 P(not F) = 1 - P(F) P(not F) = 1 - 0.17 P(not F) = 0.83

jimthompson5910 (jim_thompson5910):

A and C are correct

jimthompson5910 (jim_thompson5910):

D looks a bit odd. I don't know what to make of it

OpenStudy (anonymous):

Thank you. For D I did P(E) + P(F) - P(E and F)

jimthompson5910 (jim_thompson5910):

so they want to know P(E or F) for part D?

OpenStudy (anonymous):

Yes, sorry, typed it wrong

jimthompson5910 (jim_thompson5910):

If so, then, P(E or F) = P(E) + P(F) - P(E and F) P(E or F) = P(E) + P(F) - P(E)*P(F) P(E or F) = 0.45 + 0.17 - 0.45*0.17 P(E or F) = 0.5435

jimthompson5910 (jim_thompson5910):

P(E and F) = P(E)*P(F) only because E and F are assumed to be independent events

OpenStudy (anonymous):

thanks for your help!

jimthompson5910 (jim_thompson5910):

np

OpenStudy (paxpolaris):

d. P(E) or (F) assuming independent -> \[=P(E | \neg F)+P(F)\]

OpenStudy (paxpolaris):

=.83*.45+.17=.5435 nvm

OpenStudy (paxpolaris):

a. P(E^c) sorry, could someone explain what E^c means ... what's c?

jimthompson5910 (jim_thompson5910):

the ^c means complement a complement of a set or event is simply everything but that set or event http://en.wikipedia.org/wiki/Complement_%28set_theory%29

jimthompson5910 (jim_thompson5910):

So for example R = event that it rains that means R^c = even that it does not rain

OpenStudy (paxpolaris):

thanks :)